# MET-Bench: Multimodal Entity Tracking for Evaluating the Limitations of Vision-Language and Reasoning Models

### v1

@@ 5.1 Tracking in Text Outperforms Images
Few-shot prompting provides only marginal improvements, but chain-of-thought gives large performance increases. GPT-4o’s accuracy jumps from $36.0\%$ to $99.8\%$ in text, and from $32.4\%$ to $84.2\%$ in image tracking. These results suggest that when guided to decompose the task step-by-step, models can reason more effectively using image inputs, a finding that complements the results of performing cascaded inference in Section 5.4.
5.2 Reasoning Aids Long Sequence Accuracy
Chess
Figure 3(a) plots model accuracy against increasing sequence lengths of text actions. The models are evaluated in the zero-shot setting for sequence lengths of zero to $100$ text actions. Reasoning models like o1 and o3-mini are able to handle longer sequence lengths with a smaller decrease in accuracy. However, o1-mini performs worse than the other models as it produces more invalid FEN board representations at longer sequence lengths. In contrast, the non-reasoning models experience sharp decreases in accuracy after only a few actions. Figure 3(b) plots model accuracy against increasing sequence lengths of image actions. The models are evaluated in the zero-shot setting for sequence lengths of zero to $20$ image actions. While the reasoning models attain higher accuracies, the performance differences are smaller than in the text-action setting.
Shell Game
In the text modality in Figure 4(a), the reasoning models o1, o1-mini, and o3-mini attain the highest accuracies. o1 performs perfectly at a sequence length of 50 actions, where the non-reasoning models’ performance is significantly degraded. In the image modality results in Figure 4(b), o1 performs better than the other models, but sees a rapid decrease in accuracy with sequence lengths longer than five actions. By $20$ actions, the performance of all models has degraded to random.
The superior performance of reasoning models like o1 suggests that structured inference mechanisms, such as test-time search or chain-of-thought, are crucial for maintaining coherent entity states over long sequences. This aligns with prior findings that models trained on structured reasoning tasks (e.g., mathematics, coding) develop stronger implicit state-tracking capabilities (Kim et al., 2024), whereas standard vision-language models struggle with entity persistence beyond short contexts.
5.3 Models Understand Image Actions
(a) Shell Game zero-shot accuracy with text actions.
(b) Shell Game zero-shot accuracy with image actions.
Figure 4: In the text action setting, the reasoning models o1, o1-mini, and o3-mini achieve the highest performance over long action sequences, but performance degrades for all models as sequence length increases. In the image action setting, o1 performs the best, but achieves an accuracy no better than guessing the starting state by $20$ actions.
Model Start End Overall
Chess
GPT-4o-mini 62.1 55.2 41.0
GPT-4o 97.3 97.0 94.5
Shell Game
GPT-4o-mini 100.0 100.0 100.0
GPT-4o 100.0 100.0 100.0
Table 2: Percent image-action classification accuracy for various models. We report the accuracy of predicting the action start, end, and overall/UCI action for both Chess and Shell Game on $1{,}000$ image actions.
@@ 5.3 Models Understand Image Actions
Baseline
Game Start 74.4 33.3
Zero-Shot
GPT-4o mini 63.0 28.0
GPT-4o 89.8 34.0
Table 3: Cascaded entity tracking accuracy for Chess and Shell (Image → Text). In cascaded inference, the model is first used to map each image action to the text representation of the action. Then model is prompted to perform the entity tracking task as in the text-action setting.
Table 2 shows the performance of GPT-4o and GPT-4o-mini on classifying the text action represented by each image action. For Chess, the action ‘start’ is the square of the piece being moved, and ‘end’ is the destination square. For Shell Game, the ‘start’ is the first shell to be swapped and ‘end’ is the second. ‘Overall’ is the accuracy of classifying the entire action (start and end) correctly. While GPT-4o-mini struggles to recognize actions in chess, GPT-4o achieves an accuracy of $94.5\%$. Both models attain perfect accuracy on the simpler Shell Game domain. This indicates that perception of the image-actions is not the fundamental limiting factor for effective entity tracking with image inputs.
A potential concern is whether the observed failures stem from poor image representation rather than reasoning deficiencies. However, our cascaded inference experiments (Section 5.4) demonstrate that when models first translate image actions into text, they achieve near-text-level accuracy. This suggests that models can correctly parse image-based actions but struggle with integrating them into coherent state updates, a limitation in reasoning rather than perception.
@@ 5.4 Cascading Matches Text-Only Tracking
Table 3 shows the accuracy of cascaded inference in the zero-shot setting for GPT-4o and GPT-4o-mini. In this setting, the image actions are first translated into text actions, and then run through the text-based entity tracking pipeline. The performance in the cascaded setting is similar to the text-action performance, showing that the model has the task-knowledge needed to perform entity tracking in both domains, but cannot reason effectively in the image modality.
@@ 6 Discussion
Our evaluation of frontier model performance on MET-Bench provides several insights into the current state and remaining challenges of multimodal entity tracking. We demonstrate a significant performance gap between text-based and image-based entity tracking across all evaluated models, with even state-of-the-art vision-language-reasoning models struggling to maintain accurate entity states when processing visual inputs. This disparity persists across both the Chess and Shell Game domains, suggesting a fundamental limitation in current architectures’ ability to reason about entity states through visual observations.
This finding is particularly noteworthy given that our image-action classification results (Table 2) demonstrate that models can accurately perceive and classify individual visual actions. The gap between perception and reasoning suggests that the challenge lies not in processing visual inputs, but in maintaining and updating coherent entity information across sequential visual observations.
Our cascaded inference experiments provide further evidence for this interpretation. When models first translate visual inputs to text before performing entity tracking, they achieve performance comparable to pure text-based tracking. This indicates that the models possess the relevant task knowledge and reasoning capabilities, but struggle to apply them directly in the visual domain.
Further, the effectiveness of chain-of-thought prompting, particularly in the Shell Game domain where it improved GPT-4o’s accuracy from $36.0\%$ to $99.8\%$ for text and $32.4\%$ to $84.2\%$ for images, highlights the importance of explicit reasoning for entity tracking. This improvement indicates that current models can perform complex entity tracking when guided to decompose the task into smaller steps, even in novel domains not present in their training data. However, the fact that such prompting was necessary suggests that models do not implement robust tracking, particularly in multimodal settings. Lastly, the performance of specialized reasoning models like o1 and o3-mini on longer sequences demonstrates the potential of architectures explicitly trained for sequential reasoning to maintain coherent entity states despite the challenges of accumulating errors over extended sequences.
@@ 7 Related Work
Entity tracking has been extensively studied in textual domains, with a focus on probing and improving language models’ abilities to maintain representations of entity states. For instance, Toshniwal et al. (2022) evaluates chess as an entity tracking domain, employing fine-tuned models (Radford et al., 2019) to assess performance. Similarly, Kim and Schuster (2023) examine the impact of model size and fine-tuning on entity tracking in textual settings similar to our Shell Game domain. Tandon et al. (2020) construct a benchmark for understanding entity state changes in procedural texts. Shirai et al. (2022) construct the Visual Recipe Flow corpus and evaluate the ability of multimodal embedding models to properly sequence images depicting recipe states. In contrast, our work requires predicting entity state changes from actions specified in images and involves larger state spaces.
Several studies explore the implicit representations of entity states in language models. Li et al. (2021) and Long et al. (2016) use semantic probing to reveal that Transformer-based models (Vaswani et al., 2017) capture entity state representations implicitly during textual reasoning. Building on this, Prakash et al. (2024) demonstrate that fine-tuning language models for entity tracking tasks enhances pre-existing internal mechanisms rather than learning entirely new representations. Li et al. (2023) find that Transformers trained on Othello games form internal representations of the game state.
Efforts to improve textual entity tracking beyond domain-specific fine-tuning include Fagnou et al. (2024), which establishes theoretical limitations of the Transformer architecture in tracking entities. They propose a novel attention mechanism to enhance entity tracking in Transformers. Gupta and Durrett (2019) fine-tunes small Transformer-based models for tracking entity state in instructional texts. Kim et al. (2024) investigates how code pretraining improves language models’ abilities to track entities in text, while Yoneda et al. (2024) introduce Stalter, a prompting method designed to maintain accurate state representations in text-based robotics planning.
These works focus on entity tracking as a unimodal, text-based reasoning task. While unimodal approaches have achieved substantial progress, there remains a gap in evaluating models’ ability to integrate multimodal inputs for entity tracking. Our work extends these evaluations to the multimodal setting and quantifies the performance improvement of reasoning models for entity tracking.
@@ 8 Conclusion
Our findings suggest that the primary bottleneck in multimodal entity tracking is not visual recognition but sequential reasoning over visual updates. Unlike text-based representations, which align with the models’ training paradigms, visual updates require implicit state reconstruction—a task that current architectures do not perform reliably. Future work should explore the effect of additional visual-reasoning post-training, explicit memory structures, or hybrid symbolic representations to mitigate this gap. Additional research directions include investigating the role of entity tracking in world-modeling, narrative understanding, and expanding MET-Bench to include more complex domains beyond games. We believe addressing these challenges will be crucial for developing AI systems capable of robust reasoning for real-world tasks.
9 Limitations
Our benchmark is restricted to two synthetic domains—Chess and the Shell Game. While these domains are well-structured and offer a clear framework for assessing entity state tracking, they may not fully capture the complexity of real-world multimodal reasoning tasks that require entity tracking.
Our study highlights a substantial performance gap between text-based and image-based entity tracking, yet the exact causes of this disparity remain unclear. While our cascaded inference experiments suggest that models struggle to reason directly over visual updates rather than simply perceiving them, further investigation is needed to pinpoint the underlying source of the problem and whether it’s possible to explicitly ameliorate it.
Additionally, our fine-tuning experiments demonstrate that performance can be improved with domain-specific adaptation, but we do not investigate the trade-offs between fine-tuning and generalization across unseen multimodal datasets. Understanding whether these improvements transfer to novel multimodal reasoning tasks is an important avenue for future work.
Ethical Considerations
Vision-language models may inherit biases from their training data or other aspects of their development, leading to disparities in performance across different languages, cultures, or other categories that intersect with textual and visual representations. The evaluations in this work can be used to improve the capabilities of frontier models, which if deployed in contexts requiring multimodal entity tracking, could further expose end-users to the biases present in this work including our choice of game domains and language.
Acknowledgments
This paper and associated code were created with the assistance of ChatGPT for proofreading and debugging.
@@ A.1 Models Struggle to Integrate Mixed Modalities
A.2 Fine-Tuning Improves Multimodal Entity Tracking
Fine-tuned Text Image
Chess
GPT-4o mini 89.2 -
GPT-4o 97.0 86.4
Shell Game (S=3)
GPT-4o mini 32.0 -
GPT-4o 74.0 32.0
Table 4: Fine-tuned model entity tracking accuracy for the Chess and Shell Game domains (Text actions and Image actions).GPT-4o-mini does not support image finetuning. A training set of $100$ action sequences of length ten were used for Chess and $20$ action sequences of length three and five for Shell Game.
Fine-tuning using the OpenAI fine-tuning API substantially improves model performance across both text and image modalities, as shown in Table 4. In Chess, fine-tuned models outperform even the strongest zero-shot reasoning models, achieving 97.0% accuracy in the text domain and a significant boost to 86.4% in the image domain. This suggests that even with a relatively small dataset, fine-tuning allows the model to learn entity tracking representations that generalize better in both modalities. Notably, fine-tuning leads to a larger improvement in the image modality than in the text modality. This reinforces the idea that pretrained models already encode strong textual reasoning capabilities, whereas multimodal reasoning requires additional adaptation.
In contrast, improvements on a simplified version of Shell Game with only three moves are minimal in case of image-encoded actions. The Shell Game task is not present in the training data and it’s harder for the model to generalize, even when exposed to a large fraction of the possible games of the given length. This may indicate that the Shell Game domain is simply too challenging for the model to learn in both the image and text settings from a limited number of examples. A more complex training curriculum involving fine-tuning over multiple game lengths may be required.
Model Name Image Reasoning
Claude 3.5 Sonnet
Anthropic (2024)
Gemini-2.0-Flash
Hassabis and Kavukcuoglu (2024)
GPT-4o mini
OpenAI (2024a)
GPT-4o
OpenAI (2024b)
Minimax-VL-01
MiniMax et al. (2025)
o1-mini
OpenAI (2024c)
o1
OpenAI (2024d)
o3-mini
OpenAI (2025)
Table 5: Comparison of capabilities of language models evaluated using the MET benchmark. All evaluated models support text input and output. The total API cost of experiments run is \$$2340.00$.
Configuration Epochs LRM Batch
Chess
   Text:
    GPT-4o-mini 3 1.8 1
    GPT-4o 3 2 1
   Image:
    GPT-4o 3 2 1
Shell Game
   Text:
    GPT-4o-mini 3 1.8 1
    GPT-4o 3 2 1
   Image:
    GPT-4o 5 2 1
Table 6: Hyperparameters used for fine-tuning across domains and modalities. The training epochs, learning rate multiplier (LRM), and batch size are reported.
@@ A.2 Fine-Tuning Improves Multimodal Entity Tracking
A.3 Models
The models evaluated using MET-Bench are listed in Table 5.
Minimax-VL-01
This model is released under the license: https://github.com/MiniMax-AI/MiniMax-01/blob/main/LICENSE. The model is 465 billion parameters and is trained on a “diverse [dataset] incorporating diverse sources including academic literature, books, web content, and programming code” and post-training dataset encompassing many multimodal and NLP tasks of 512 billion tokens (MiniMax et al., 2025).
A.3.1 Proprietary Models
These models have limited information about their training and development. Like Minimax-VL-01, these models are likely trained on diverse, web-scale corpora spanning many domains and tasks. We provide links to the current terms of their use.
Claude 3.5 Sonnet
Gemini-2.0-Flash
GPT-4o mini, GPT-4o, o1, o1-mini, o3-mini
A.4 Datasets
The Chess dataset is adapted from Toshniwal et al. (2022) which is adapted from the MillionBase dataset, available for download at https://rebel13.nl/rebel13/rebel%2013.html. To the best of our knowledge, no license or terms of use are currently listed for either the original MillionBase dataset or dataset of Toshniwal et al. (2022). Our usage of this dataset is consistent with the description of its use by Toshniwal et al. (2022).
MET-Bench is intended for evaluating and improving the ability of VLMs to perform entity tracking. It is released under the MIT License.

### v2

@@ 4.1 Tracking in Text Outperforms Images
Chain-of-thought and reasoning lead to large performance increases and multiple models attain near perfect accuracy in text. These results suggest that when guided to decompose the task step-by-step, models can reason more effectively (but still largely imperfectly) using image inputs, a finding that complements the results of performing cascaded inference as discussed below.
Table 3: Cascaded entity tracking accuracy for Chess and Shell (Image → Text). In cascaded inference, the model is first used to map each image action to the text representation of the action. The model is prompted to perform the entity tracking task as in the text-action setting.
@@
Chain-of-Thought
Gemma3 27B 55.7 63.0
GPT-4o mini 66.4 47.6
GPT-4o 93.2 99.4
4.2 Reasoning Aids Long Sequence Accuracy
We varied the sequence length of the actions to quantify the effect of compounding errors on the models. These sequences range from zero to $100$ actions in both the image and text action modality. This serves to quantify the relative drop-off in performance among models in the reasoning and chain-of-thought settings.
Chess
Figure 3(a) plots model accuracy against increasing sequence lengths of text actions. The models are evaluated in the chain-of-thought setting for sequence lengths of zero to $100$ text actions. Reasoning models like Claude 3.7 Sonnet Thinking are able to handle longer sequence lengths with a smaller decrease in accuracy. In contrast, the non-reasoning models experience sharp decreases in accuracy after only a few actions. Figure 3(b) plots model accuracy against increasing sequence lengths of image actions. The models are evaluated in the chain-of-thought setting for sequence lengths of zero to $20$ image actions.
Shell Game
In the text modality in Figure 4(a), the reasoning models Gemini 2.5 Pro and Claude 3.7 Sonnet Thinking attain the highest accuracies. Gemini 2.5 Pro performs nearly perfectly at a sequence length of $100$ actions, where the non-reasoning models’ performance is significantly degraded. In the image modality results in Figure 4(b), Gemini 2.5 Pro performs better than the other models, which see a rapid decrease in accuracy with sequence lengths longer than ten actions. By $20$ actions, the performance of all models except Gemini 2.5 Pro has degraded to random guessing.
The superior performance of reasoning models suggests that structured inference mechanisms, such as test-time search or chain-of-thought, are crucial for maintaining coherent entity states over long sequences. This aligns with prior findings that models trained on structured reasoning tasks (e.g., mathematics, coding) develop stronger implicit state-tracking capabilities (kim2024codepretrainingimprovesentity), whereas standard vision-language models struggle to track entities beyond short visual action sequences.
Table 4: Percent image‐action classification accuracy (with 95% CI) for various models. We report the accuracy of predicting the shell swap/UCI action for both Chess and Shell Game on 1,0001{,}000 image actions.
Model Perception Accuracy (%)
Chess
GPT-4o-mini 46.4 ± 1.0
GPT-4o 94.5 ± 1.4
Claude 4.5 Opus 100.0 ± 0.2
Gemma3 27B 67.5 ± 2.9
Shell Game
GPT-4o-mini 100.0 ± 0.2
GPT-4o 100.0 ± 0.2
Claude 4.5 Opus 100.0 ± 0.2
Gemma3 27B 100.0 ± 0.2
4.3 Models Understand Image Actions
Table 4 shows the performance of models on predicting the text action represented by each image. While some models fail to accurately localize the Chess moves, the evaluated models understand the Shell Game actions with high accuracy. This indicates that perception of the image-actions is not the only limiting factor for effective entity tracking with image inputs.
@@ 4.4 Cascading Matches Text-Only Tracking
To differentiate the effects of perception and visual reasoning, we evaluate cascaded inference. The performance in this cascaded setting is similar to the text-action performance, showing that the model has the task-knowledge needed to perform entity tracking in both domains, but cannot reason effectively directly from the image modality. Using the text actions predicted from the images in the image-action classification task, we devise a cascaded approach to test the effect of performing multimodal inference in two separate steps. Given a sequence of image-based observations $\bm{a_{t}^{\text{image}}}$, a vision-language model (VLM) first predicts the corresponding text-based action sequence $\hat{\bm{a}}^{\text{text}}=g(\bm{a}^{\text{image}})$, where $g$ maps images to text using the procedure described in the previous section. The text actions are then used for zero-shot inference to estimate the final state as $\bm{S_{T}}=f(\bm{S_{0}},\hat{\bm{a}}^{\text{text}})$. This removes the need for the model to perform entity tracking directly in the image modality, isolating the effect of perceptual failures. Table 3 shows results for this approach in the zero-shot setting. The image actions are first explicitly translated into text actions, and then run through the text-based entity tracking pipeline.
4.5 GRPO Improves Entity Tracking
We train open-source VLMs for improving entity tracking on a 5-action version of MET-Bench. GRPO is highly effective at teaching entity tracking to models, achieving performance on par with several advanced closed models. However, the lack of cross-modal transfer suggests models process entity tracking in text and image through different pathways. Robust multimodal entity tracking may require architectural innovations or training strategies beyond what RL provides.
Table 5: Entity tracking accuracy (%) after RL training for Gemma3 4B and 27B models. GPT-4o-mini model for comparison. 95% confidence intervals.
Model Text Image
Shell Game
Gemma3-4B (Base) 44.4 $\pm$ 4.3 32.2 $\pm$ 4.1
Gemma3-4B (GRPO Text) 71.0 $\pm$ 4.0 32.2 $\pm$ 4.1
Gemma3-27B (Base) 80.4 $\pm$ 3.5 44.4 $\pm$ 4.3
Gemma3-27B (GRPO Image) 79.8 $\pm$ 3.5 62.0 $\pm$ 4.2
GPT-4o-mini 86.4 $\pm$ 3.0 36.8 $\pm$ 4.2
Chess
Gemma3-4B (Base) 66.1 $\pm$ 1.2 14.6 $\pm$ 0.9
Gemma3-4B (GRPO Text) 84.4 $\pm$ 0.9 26.3 $\pm$ 1.1
Gemma3-27B (Base) 75.8 $\pm$ 1.0 63.1 $\pm$ 1.2
Gemma3-27B (GRPO Image) 73.8 $\pm$ 1.1 74.5 $\pm$ 1.1
GPT-4o-mini 77.0 $\pm$ 0.5 60.1 $\pm$ 0.5
Table 5 shows the results of GRPO training on the Shell Game domain. When trained on text-only inputs with chain-of-thought prompting, Gemma3-4B improves from 44.4% (near random) to 71.0% accuracy. Figure 5 illustrates the learning dynamics.
Limited Cross-Modal Transfer
Despite the improvement in single modality entity tracking, RL reasoning training does not transfer robustly to the held-out modality. In Shell Game, the held-out modality does not improve at all. This finding suggests that the entity tracking capabilities learned through text-based RL training are modality-specific and do not generalize to other modalities. The apparent transfer in Chess image modality can be attributed to the model learning to output more valid FENs.
Shell, Image Actions (Gemma3 27B) Shell, Text Actions (Gemma3 4B)
Refer to caption Refer to caption
Chess, Image Actions (Gemma3 27B) Chess, Text Actions (Gemma3 4B)
Refer to caption Refer to caption
Figure 5: GRPO training curves across domains and modalities. Each series shows test accuracy for the train modality and transfer modality for image ($\bullet$) and text ($\blacksquare$). GRPO improves in-modality performance but this does not substantially transfer to the other modality. 95% confidence intervals.
@@ 5 Discussion
Many VLM applications rely on robust multimodal entity tracking. Our evaluation of frontier models on MET-Bench provides several insights into the current remaining challenges of multimodal entity tracking. We demonstrate a significant performance gap between text-based and image-based entity tracking across many state-of-the-art vision-language-reasoning models. This disparity generally persists across both the Chess and Shell Game domains, suggesting a fundamental limitation in current architectures’ ability to reason about entity states through visual observations.
This finding is particularly noteworthy given that our image-action classification results (Table 4) demonstrate that models can accurately perceive and classify individual visual actions in Shell Game. The gap between perception and reasoning suggests that the challenge lies not in processing visual inputs, but in maintaining and updating coherent entity information across sequential visual observations.
Our cascaded inference experiments provide further evidence for this interpretation. In Shell Game, when models first translate visual inputs to text before performing entity tracking, they achieve performance comparable to pure text-based tracking. This indicates that the models possess the relevant task knowledge and reasoning capabilities, but struggle to apply them directly in the visual domain.
Further, the effectiveness of chain-of-thought prompting, particularly in the Shell Game domain where it improved Claude 3.7 Sonnet’s accuracy from near random to $100.0\%$ for text and $77.4\%$ for images, highlights the importance of explicit reasoning for entity tracking. Current models can perform complex entity tracking when guided to decompose the task into smaller steps, even in novel domains not explicitly present in their training data.
The RL experiments demonstrate that reasoning training can substantially improve entity tracking within a single modality, with Gemma3-4B improving from near-random performance to 71.0% accuracy on text-based Shell Game tracking. However, the lack of robust cross-modal transfer suggests that current VLM architectures learn modality-specific representations rather than unified entity tracking capabilities. Achieving robust multimodal entity tracking may require new methods, such as explicit cross-modal alignment objectives or other inductive biases that encourage models to develop modality-agnostic tracking mechanisms.
@@ 6 Related Work
Entity tracking has been extensively studied in textual domains, with a focus on probing and improving LLMs’ abilities to maintain representations of entity states. toshniwal2022chess evaluates chess as an entity tracking domain, employing finetuned models (radford2019language) to assess performance. Kim and Schuster (2023) examine the impact of model size and fine-tuning on entity tracking in textual settings similar to our Shell Game domain. Tandon et al. (2020) construct a benchmark for understanding entity state changes in procedural texts. Shirai et al. (2022) construct the Visual Recipe Flow corpus and evaluate the ability of multimodal embedding models to properly sequence images depicting recipe states. In contrast, our work requires predicting entity state changes from actions specified in images and involves larger state spaces.
Several studies explore the implicit representations of entity states in LLMs. li2025language employ interpretability methods to understand how LLMs implement entity tracking. Li et al. (2021) and Long et al. (2016) use semantic probing to reveal that Transformer-based models (vaswani2017attention) capture entity state representations implicitly during textual reasoning. Building on this, prakash2023fine demonstrate that fine-tuning language models for entity tracking tasks enhances pre-existing internal mechanisms rather than learning entirely new representations. li2023emergent find that Transformers trained on Othello games form internal representations of the game state.
Efforts to improve textual entity tracking beyond domain-specific fine-tuning include Fagnou et al. (2024), which establishes theoretical limitations of the Transformer architecture in tracking entities. They propose a novel attention mechanism to enhance entity tracking in Transformers. Gupta and Durrett (2019) trains small Transformer-based models for tracking entity state in instructional texts. kim2024codepretrainingimprovesentity investigates how code pretraining improves language models’ abilities to track entities in text, while yoneda2024statler introduce Statler, a prompting method designed to maintain accurate state representations in text-based robotics planning.
Recent work demonstrates the effectiveness of reinforcement learning for improving reasoning capabilities in language models. DeepSeek-R1 (deepseekai2025deepseekr1incentivizingreasoningcapability) shows that RL training with verifiable rewards can significantly enhance mathematical and logical reasoning with Group Relative Policy Optimization (GRPO) (shao2024deepseekmath). We apply GRPO to multimodal entity tracking, demonstrating both its promise for teaching visual reasoning about state updates and its limitations in enabling cross-modal transfer.
@@ 7 Conclusion
Our findings suggest that the primary bottleneck in multimodal entity tracking may not be visual recognition but sequential reasoning over visual updates. Unlike text-based representations, which align with the models’ training paradigms, visual updates require implicit state reconstruction, a task that most current architectures do not perform reliably.
Our RL experiments reveal a promising path forward that can effectively train open-source models, achieving performance comparable to closed models. However, the lack of cross-modal transfer underscores that visual entity tracking requires fundamentally different representations or training objectives. Future work should explore deficits in entity tracking using mechanistic interpretability methods. Additional research directions include novel multimodal representation learning methods and explicit memory structures to mitigate this gap. We believe addressing these challenges will be crucial for developing AI systems capable of robust multimodal reasoning for real-world tasks.
Impact Statement
As with any work that improves AI reasoning capabilities, advancements in multimodal entity tracking could eventually contribute to more capable autonomous systems. While our benchmark focuses on structured game domains (Chess and Shell Game), the underlying capabilities are relevant to robotics and other systems employing automated decision-making. We encourage researchers building on this work to consider appropriate safeguards in consequential real-world settings.
@@ A.1 Models Struggle to Integrate Mixed Modalities
A.2 GRPO Training Hyperparameters
Table 6 summarizes the key hyperparameters used for GRPO training across domains and modalities. All experiments are implemented using the TRL library (vonwerra2020trl) with the AdamW optimizer. For the 27B model, we employ LoRA (hu2021loralowrankadaptationlarge) with rank 16 to reduce memory requirements. All experiments use chain-of-thought prompting and are trained on 100 examples with 5 actions per sequence.
Table 6: GRPO training hyperparameters for the experiments shown in Figure 5.
Gemma3-4B Gemma3-27B
Param Chess Shell Chess Shell
(Text) (Text) (Image) (Image)
Optimizer AdamW AdamW AdamW AdamW
Learning Rate 1e-6 1e-6 1e-5 1e-5
Batch Size 4 4 1 1
Grad. Accum. 16 16 16 16
Group Size 8 8 8 8
LoRA r=16 r=16